Solve and check
3/x-1 + 2/y+2 - 1/z+1 =9
1/x-1 - 4/y+2 + 3/z+1 =-5
2/x-1 - 2/y+2 - 5/z+1 =7

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

Solve and check
3/x-1 + 2/y+2 - 1/z+1 =9
1/x-1 - 4/y+2 + 3/z+1 =-5
2/x-1 - 2/y+2 - 5/z+1 =7
--------------
Sub r, s & t for 1/(x-1), 1/(y=2), etc.
===============
3r + 2s - t = 9
r - 4s + 3t = -5
2r - 2s - 5t = 7
---
Solve for r, s & t
x = (1/r) + 1
etc.

Answer by ikleyn(52937)

.
For more such problems see the lesson
    - Solving systems of non-linear equations in three unknowns using Cramer's rule
in this site.